OpenStudy (anonymous):

how do you solve this problem using the zero product property: (4x +5)(x + 7) = 0

OpenStudy (anonymous):

If: \[(x-a)(x-b)=0; a,b \ge 0 \implies x=a,x=b\]

OpenStudy (anonymous):

So: \[(4x+5)(x+7)=0 \implies 4x=5\implies x=\frac{5}{4}, x=-7\]

OpenStudy (anonymous):

Because: \[(0)(x)=0 \] For any x so only one of the factors need be zero.